# A Comprehensive Survey on Automatic Text Summarization with Exploration of LLM-Based Methods

### v2

@@ 2.3 Related Surveys on ATS
Several ATS surveys have been published over the years, generally adopting a technical categorization approach by classifying ATS methods as either “extractive” or “abstractive”, which are summarizaed in Table 1. For instance, studies like [79, 80] focus on “extractive” methods, which select key sentences or paragraphs from the original documents to form concise summaries. These surveys cover techniques such as term frequency, statistical models, and supervised learning approaches. Similarly, [12] highlights extractive methods that rely on neural models. In contrast, research like [13, 14] explores abstractive summarization methods, documenting the transition from early statistical and rule-based techniques to the latest advancements in neural language models. Additionally, [76] provides a comprehensive review of both extractive and abstractive approaches.
@@ 3.1 Definitions for Automatic Text Summarization
Figure 1: ATS Framework: Detailed Process from Data Acquisition through Text Pre-processing, Summarization Modeling to Evaluation Metrics.
@@ 3.2 Process of Automatic Text Summarization
    Summarization Modeling. Modeling, the cornerstone of an ATS system, is dedicated to developing versatile language models that can interpret and distill language data into concise summaries. This is achieved through rule-based, statistical, or deep learning approaches which extract patterns from the language data. The process of modeling in ATS is inherently an NLP task, typically commencing with language modeling and subsequently progressing to summarization modeling. The conventional ATS methods are bifurcated into language modeling and summarization modeling. For an in-depth discussion and categorization of the various ATS models, refer to Section 6. This section also provides a detailed exploration of methods based on Large Language Models (LLMs).
@@
3.3 Methodology for Crawling ATS Papers
We developed an automated crawling algorithm (Algorithm 1) to efficiently collect papers relevant to ATS, addressing the limitations of manual keyword-based searches. By integrating web crawling and Large Language Model (LLM)-based filtering, the algorithm enhances both the accuracy and comprehensiveness of the retrieval process.
To this end, we designed a three-stage algorithm that automates the collection, filtering, and categorization of ATS papers. The process leverages synonym-enhanced keyword searches and analyzes both titles and abstracts for relevance, which are as following:
  1. 1.
  2. 2.
    Relevance Filtering: Retrieve abstracts to filter out irrelevant papers. Since Google Scholar often provides incomplete abstracts, the algorithm additionally retrieves their publication URLs ($pub\_url$) to scrape full abstracts. For URLs pointing to PDFs, text is extracted directly from the files using PyPDF.
  3. 3.
@@ 3.3 Methodology for Crawling ATS Papers
Initialize the LLM for classification $LLM$
Define search keywords $K=\{\texttt{summarization},\texttt{text condensation},\dots\}$
@@
     $results\leftarrow C.search(k)$triangleright Perform search using keyword $k$
@@
     if $paper.pub\_url$ ends with .pdf then
         Extract $abstract$ from $page.FILE$
@@
         Extract $abstract$ from $page.HTML$
@@
     Clean and store $abstract$ in $D$
@@
     Add $result$ to categorized dataset $R$
@@ 4 ATS Datasets
Open-source datasets are mostly adopted for general ATS tasks but often lack suitability for domain-specific applications, as noted in studies [99, 100, 101]. To address these gaps, we outline methodologies for creating newly custom datasets, focusing on leveraging Large Language Models (LLMs). LLMs enable scalable and efficient generation of domain-specific data, reducing manual effort and enhancing dataset diversity.
@@ 4.1 Open-source Datasets
We collected datasets in the ATS domain that are publicly accessible online or via email, referred to as Open-Source Datasets. We adopted the automated retrieval algorithm proposed in Sec. 3.3 to retrieve the ATS datasets, by using ATS dataset related queries. This process also consists of three main steps:
@@
    Relevance Filtering: To exclude irrelevant articles, we employed the LLaMa3 LLM for few-shot learning classification. Titles and abstracts were analyzed and categorized into two groups: articles introducing new datasets, labeled as “dataset,” and those unrelated, labeled as “non-dataset.” Manual screening was then conducted to ensure the accuracy of the classification, resulting in a definitive list of relevant datasets.
@@
    Core Datasets: These are datasets with over 100 citations, considered central to the field. Detailed introductions for each core dataset are provided, and their key information is summarized in Table 2. To ensure completeness, we reviewed related surveys and integrated any significant datasets they mentioned that were not retrieved through our search.
@@
    Supplementary Datasets: These are datasets with fewer than 100 citations, providing additional resources for specific or emerging use cases. Their basic information is summarized in Table 3.
Table 2: Core datasets with more than 100 citations in the ATS. The table is arranged in ascending order by publication year and provides details on the publish year, size (number of total pairs including train, validation, and test), domain, language, along with their publicly available URLs for download.
@@
4.1.1 Core Datasets
The table of core datasets includes details on the publish year, size (number of total pairs including train, validation and test), domain, language, along with their public urls for download. In the following segments, we overview the datasets listed in Table 2. The following is a brief introduction to each of the core datasets.
@@ 4.1.1 Core Datasets
LCSTS[105] contains over 2 million Chinese micro-blogs from domains inlcude politics, economics, and entertainment, sourced from Sina Weibo’s official accounts such as People’s Daily, the Ministry of National Defense, etc.. Summaries are manually annotated, with rules applied to ensure quality (e.g., accounts with over 1 million followers). The dataset is divided into three parts: a large master dataset, a high-quality subset with manual scoring, and a refined test set.
@@
XL-sum[117] is a multilingual abstractive summarization dataset with 1 million article-summary pairs in 44 languages, sourced from BBC. It focuses on diverse and low-resource languages, offering high-quality, concise summaries. Fine-tuned mT5 models achieve competitive results, making XL-sum the largest dataset for multilingual summarization in both samples and language coverage.
@@
4.1.2 Supplementary Datasets
Supplementary datasets refer to those cited fewer than 100 times. Although less frequently used than conventional datasets, they offer unique diversity and can serve as valuable resources for addressing specific challenges in the ATS task. The information is presented in Table 3.
Table 3: Supplementary datasets with fewer than 100 citations in ATS. The table is formatted in the same manner as the core dataset table, exhibiting diversity and serving as a valuable complement to ATS datasets.
@@ 4.1.2 Supplementary Datasets
[146] MM-AVS 2021 2,260 Multimodal English https://github.com/xiyan524/MM-AVS
@@
As summarized in Tables 2 and 3, current summarization datasets offer several strengths: they include large-scale datasets for training deep neural networks and smaller, refined datasets for evaluation, are predominantly open source and easily accessible, and primarily focus on the news domain. However, there is a gap in high-quality datasets for other domains, limiting research and practical applications in specialized areas such as financial earnings releases [76]. To address this need, we propose methodologies for constructing customized datasets, which are detailed in the following section.
@@ 4.2 Techniques to Build New Datasets
LLM-based Data Generation: LLMs are effective for generating summaries. A straightforward approach is to input original texts along with carefully designed prompts, then fine-tune the model to better grasp the summarization task and produce high-quality outputs (see Section 6.2). Some studies focus on LLM-based annotation methods. For example, [165] used DialoGPT[166] to transform annotation into tasks like keyword extraction and redundancy detection, improving informativeness and relevance. [167] developed a GPT-3 algorithm for few-shot annotation of medical dialogues, outperforming human-trained models in accuracy and coherence. LLM-based data generation could also be incorporated with human annotators to further improve the robustness. For instance, [168] proposed a GPT-3-based scheme combining LLM generation with human evaluation for improved outcomes.
@@ 5.1 Pre-processing methods
Stemming: is the process of converting words with the same root or stem to a basic form by eliminating variable endings like ”es” and ”ed” [177, 178]. This process was initially employed in open vocabulary text mining, for similar reasons as removing stop words: to reduce computation time and enhance recall in information retrieval [179]. [180] proposed reducing each word to its initial letters, a method referred to as Ultra-stemming, which demonstrated significant improvements.
Sentence Segmentation: splits texts into sentences. The simplest method involves using end markers such as ”.”, ”?”, or ”!”, but it is prone to a lot of interference, including abbreviations like ”e.g.” or ”i.e.”. To address this issue, simple heuristics and regular expressions are employed [181]. Sentence segmentation is typically necessary for processing long texts [182] or for specific model designs, such as using sentences as nodes in a graph [183].
@@ 5.2 Pre-process Toolkits
The primary tools for English pre-processing with Python are NLTK [^1] https://www.nltk.org/ and TextBlob [^2] https://github.com/sloria/TextBlob. NLTK (Natural Language Toolkit) stands out as a leading platform for developing Python programs that handle human language data. It offers a suite of text processing libraries covering classification, tokenization, stemming, tagging, parsing, and semantic reasoning. NLTK also provides wrappers for industrial-strength NLP libraries and maintains an active discussion forum. TextBlob, built on NLTK, offers a simplified API for various NLP tasks, including part-of-speech tagging, noun phrase extraction, sentiment analysis, classification, translation, and more. While TextBlob is application-oriented and user-friendly, it sacrifices some flexibility compared to NLTK. For other languages, Jieba [^3] https://github.com/LiveMirror/jieba and HanLP [^4] https://github.com/hankcs/HanLP provide effective solutions for Chinese language. Jieba is a relatively lightweight tool with its primary function being word segmentation, capable of fulfilling most Chinese word splitting needs. On the other hand, HanLP[186] serves as an NLP toolkit built on PyTorch and TensorFlow, contributing to advancing state-of-the-art deep learning techniques in both academia and industry. It has demonstrated notable results in entity segmentation.
@@
6 Overview of Summarization Models
We classify summarization methods into two categories: Conventional ATS Methods, which encompass traditional approaches like extractive and abstractive models. These methods are typically limited to a specific generation paradigm, i.e., either extractive or abstractive. In contrast, with the advancement of large language models (LLMs), we collect the ATS paper based LLM, as the category of LLM-based Methods, which are more flexible and not confined to a pre-designed generative paradigm.
@@ 6 Overview of Summarization Models
Extractive summarization Statistical-based Low resource usage, simple implementation. Prone to irrelevant text and high-scoring unimportant sentences. [187, 188, 189]
@@
Abstractive summarization Tree-based Creates organized, concise summaries. May miss contextual relationships in long contexts. [199, 200]
@@
Hybrid summarization E2SA Enhances readability, preserves key information High complexity, computation cost. [211, 212, 213]
@@
6.1 Conventional ATS Methods
As shown in section 2.2, conventional ATS methods are typically classified into extractive, abstractive, and hybrid approaches. Figure LABEL:fig:classify_fine provides a detailed categorization of these methods.
6.1.1 Extractive Summarization
Extractive summarization selects sentences from the original text through three main steps: calculating sentence importance, sorting sentences based on their importance, and selecting the top-k sentences to form the summary. Sentence importance can be calculated using unsupervised or supervised methods. Extractive ATS based on unsupervised approaches rely on algorithms such as statistical, clustering, and topic-based techniques, offering high performance with low resource requirements, which will be introduced in detail below:

### v3

@@ 2.3 Related Surveys on ATS
Several ATS surveys have been published over the years, generally adopting a technical categorization approach by classifying ATS methods as either “extractive” or “abstractive”, which are summarized in Table 1. For instance, studies like [72, 73] focus on “extractive” methods, which select key sentences or paragraphs from the original documents to form concise summaries. These surveys cover techniques such as term frequency, statistical models, and supervised learning approaches. Similarly, [12] highlights extractive methods that rely on neural models. In contrast, research like [13, 14] explores abstractive summarization methods, documenting the transition from early statistical and rule-based techniques to the latest advancements in neural language models. Additionally, [76] provides a comprehensive review of both extractive and abstractive approaches.
@@ 3.1 Definitions for Automatic Text Summarization
Figure 1: Overview of the ATS (Automatic Text Summarization) framework, comparing conventional and LLM-based methods. Traditional approaches follow a staged pipeline, including data acquisition, pre-processing, modeling, and evaluation, using rule-based techniques and NLP tools. In contrast, LLM-based methods streamline the process via end-to-end summarization without explicit pre-processing.
@@ 3.2 Process of Automatic Text Summarization
    Summarization Modeling. Modeling, the cornerstone of an ATS system, is dedicated to developing versatile language models that can interpret and distill language data into concise summaries. This is achieved through rule-based, statistical, or deep learning approaches which extract patterns from the language data. The process of modeling in ATS is inherently an NLP task, typically commencing with language modeling and subsequently progressing to summarization modeling. The conventional ATS methods are bifurcated into language modeling and summarization modeling. For an in-depth discussion and categorization of the various ATS models, refer to Section 6 for conventional approaches and Section 7 for methods based on Large Language Models (LLMs).
@@
Based on the framework, the core distinction lies in the autonomous closed-loop capability of LLM for ATS, which fundamentally differs from traditional methods. Conventional ATS systems rely on fragmented pipepline regquried manual interventions, such as preprocessing of the input texts and sequential stages of language modeling followed by summarization modeling; whereas LLMs operate as unified end-to-end systems that bypass these intermediate steps entirely, as illustrated in Figure 1. This is evident in the visual contrast between the top (conventional) and bottom (LLM-based) halves of the figure: LLMs directly process raw text inputs through self-supervised learning mechanisms, eliminating the need for explicit text normalization or modular separation. Furthermore, LLMs inherently support self-contained data annotation via techniques like prompt engineering and retrieval-augmented generation, enabling dynamic synthesis of training data. The evaluation phase is similarly integrated, as LLMs can self-assess generated summaries through internal consistency checks or adversarial sampling, forming a closed-loop optimization system. This holistic capability allows LLMs to autonomously traverse data construction, modeling, and evaluation phases with minimal external intervention.
Based on the framework, the core distinction is the autonomous closed‑loop capability of LLMs, which fundamentally differs from traditional methods. Conventional ATS systems rely on fragmented pipelines requiring manual interventions, i.e., preprocessing input texts and training separate language‑modeling and summarization stages; whereas LLMs operate as unified end‑to‑end systems that bypass these steps entirely (Figure 1). This contrast is evident in the visual comparison of the top (conventional) and bottom (LLM‑based) sections: LLMs directly process raw text via self‑supervised learning, eliminating the need for explicit text normalization and modular separation. Moreover, LLMs inherently support self‑contained data annotation through techniques like prompt engineering and retrieval‑augmented generation, enabling dynamic synthesis of training data. The evaluation phase is likewise integrated—LLMs can self‑assess generated summaries via internal consistency checks or adversarial sampling, completing a closed‑loop optimization. This holistic capability enables LLMs to traverse data collection, modeling, and evaluation autonomously with minimal external intervention.
However, LLM-based methods also present challenges: while they generate fluent, contextually rich summaries, they occasionally produce factually inconsistent outputs compared to traditional statistical models. These trade-offs are discussed in detail in Section 7 alongside mitigation strategies like Retrieval Augmented Generation for ATS.
@@
Initialize the LLM for classification LLMLLM
Define search keywords $K=\{\texttt{summarization},\dots\}$
@@
  $results\leftarrow C.search(k)$ $\triangleright$ Perform search using keyword $k$
@@
  if paper.pub_urlpaper.pub\_url ends with .pdf then
   Extract abstractabstract from $page.FILE$
@@
   Extract abstractabstract from $page.HTML$
@@
  Clean and store abstractabstract in $D$
@@
  Add resultresult to categorized dataset $R$
@@
3.3 Methodology for Crawling ATS Papers
We programmed an automated crawling algorithm (Algorithm 1) to efficiently collect papers relevant to ATS, addressing the limitations of manual keyword-based searches. By integrating web crawling and Large Language Model (LLM)-based filtering, the algorithm enhances both the accuracy and comprehensiveness of the retrieval process.
To this end, we designed a three-stage algorithm that automates the collection, filtering, and categorization of ATS papers. The process leverages synonym-enhanced keyword searches and analyzes both titles and abstracts for relevance, which are as follows:
  1. 1.
  2. 2.
    Relevance Filtering: Retrieve abstracts to filter out irrelevant papers. Since Google Scholar often provides incomplete abstracts, the algorithm additionally retrieves their publication URLs (pub_urlpub\_url) to scrape full abstracts. For URLs pointing to PDFs, text is extracted directly from the files using PyMuPDF [^1] https://github.com/pymupdf/PyMuPDF.
  3. 3.
Table 2: Evaluation Results of the Proposed Retrieval Algorithm for LLM-Based ATS Papers
Task Task Description Accuracy Macro-F1 Micro-F1
is_LLM Determine whether the article discusses LLM-based ATS methods. 94.2 N/A N/A
is_dataset Identify whether the article introduces or utilizes a dataset for ATS tasks. 93.4 N/A N/A
Type Classify the article’s contribution type (e.g., survey, novel method, evaluation). 85.6 71.9 85.6
Methodology Identify the methodological approach used (e.g., prompting, fine-tuning, distillation). 86.0 85.9 86.1
Domain Determine the application domain of the article; mark as general if unspecified. 92.0 83.6 92.0
3.4 Evaluation Methodology and Results of LLM-based ATS Retrieval Algorithm 1
To evaluate the effectiveness of the retrieval algorithm in identifying ATS-related papers, we manually annotated 1,405 papers retrieved using Algorithm 1. Each paper was labeled across five dimensions: whether it involves LLM-based ATS methods (is_LLM), whether it introduces or uses a dataset for ATS tasks (is_dataset), the type of contribution (e.g., survey, novel method, evaluation; Type), the employed methodology (e.g., prompting, fine-tuning; Methodology), and the application domain within ATS (Domain). For each classification task, we report Accuracy, Macro-F1, and Micro-F1 scores. A summary of the evaluation results is presented in Table 2.
To establish ground-truth labels for evaluation, we employed a structured manual annotation protocol. Three PhD students with research expertise in natural language processing and data mining were recruited to independently annotate each of the 1,405 papers retrieved by Algorithm 1. The annotation covered five dimensions as outlined previously.
To ensure labeling quality and reduce individual bias, we adopted a majority voting strategy: for each instance, the final label was determined if at least two of the three annotators agreed. Disagreements without a majority consensus were discussed and resolved through joint review. All annotators received detailed annotation guidelines and conducted a pilot annotation round to calibrate their understanding before the full-scale task. For more detailed information, please refer to A.
3.4.1 Retrieval Algorithm Result Analysis
The evaluation results, summarized in Table 2, demonstrate that the proposed retrieval algorithm enables highly effective classification across five annotation dimensions. The binary classification tasks (is_LLM and is_dataset) achieved strong performance, with accuracy scores of 94.2% and 93.4%, respectively, indicating high precision in identifying whether articles pertain to LLM-based methods or ATS datasets.
For the multi-class tasks, the model also performed robustly. The Type task achieved an accuracy of 85.6%, with a macro-F1 score of 71.9 and a micro-F1 of 85.6, suggesting that while the model performed well on common categories, there was room for improvement in handling less frequent types. The Methodology task yielded a macro-F1 of 85.9 and a micro-F1 of 86.1, reflecting consistent performance across both dominant and rare techniques. The Domain classification task also showed strong results, with 92.0% accuracy, 83.6 macro-F1, and 90.0 micro-F1, confirming the model’s ability to distinguish between general and domain-specific ATS applications.
3.4.2 Error Analysis and Implications of Retrieval Algorithm
Two primary sources of classification error were identified:
  1. 1.
    Terminological Ambiguity: Classification performance, particularly for the is_LLM and Methodology dimensions, was affected by implicit or ambiguous phrasing in the original texts. For example, references to models such as “Transformer” or “pretrained language model,” without explicitly identifying them as LLMs, led to misclassification. Additionally, overlapping or loosely defined methodological terms hindered accurate and consistent categorization.
  2. 2.
    Long-Tail Distribution and Class Imbalance: In multi-class tasks such as Type and Methodology, a small number of dominant categories (e.g., model-centric studies) appeared much more frequently than others. This skewed distribution contributed to performance disparities, as reflected in the divergence between macro-F1 and micro-F1 scores. While high-frequency classes inflated the micro-F1, the macro-F1 revealed reduced effectiveness in identifying less-represented categories.
Despite the two errors, the overall annotation and classification framework demonstrated strong performance across all dimensions, indicating the reliability of the retrieval algorithm for ATS-related literature. In addition, the algorithm shows potential generalizability, which could potentially be expanded to other types of papers collection by altering the prompting words. Future improvements may include integrating context-aware disambiguation mechanisms, leveraging domain-specific lexicons, and applying data augmentation techniques to mitigate long-tail effects and enhance the representation of underrepresented categories.
@@ 4 ATS Datasets
Open-source datasets are mostly adopted for general ATS tasks but often lack suitability for domain-specific applications, as noted in studies [97, 98, 99]. To address these gaps, we outline methodologies for creating new custom datasets, focusing on leveraging Large Language Models (LLMs). LLMs enable scalable and efficient generation of domain-specific data, reducing manual effort and enhancing dataset diversity.
@@ 4.1 Open-source Datasets
We collected datasets in the ATS domain that are publicly accessible online or via email, referred to as Open-Source Datasets. We adopted the automated retrieval algorithm illustrated in Sec. 3.3 to retrieve the ATS datasets by using ATS dataset-related queries. This process also consists of three main steps:
@@
    Relevance Filtering: To exclude irrelevant articles, we employed an LLM for few-shot learning classification. Titles and abstracts were analyzed and categorized into two groups: articles introducing new datasets, labeled as “dataset,” and those unrelated, labeled as “non-dataset”. Manual screening was then conducted to ensure the accuracy of the classification, resulting in a definitive list of relevant datasets.
@@
    Conventional Datasets: These are well-established datasets that have been widely adopted in ATS research and have each received over 100 citations. They are considered foundational benchmarks for model training and evaluation. For each dataset, we provide a brief contextual overview in the text and summarize key attributes, such as publication year, size, domain, language, and accessibility, in Table 3.
@@
    Supplementary and Emerging Datasets: This category includes datasets with fewer than 100 citations, consisting of both recently released and domain-specific resources that are still gaining traction. While these datasets are less established, which are summarized in Table 4.
Table 3: Conventional Datasets with more than 100 citations in the ATS. The table is arranged in ascending order by publication year and provides details on the publish year, size (number of total pairs including train, validation, and test), domain, language, along with their publicly available URLs for download.
@@
4.1.1 Conventional Datasets for ATS
This subsection presents a collection of widely adopted and historically influential datasets commonly used in Automatic Text Summarization (ATS) research. These datasets, many of which have accumulated over 100 citations, serve as conventional benchmarks for training and evaluating summarization models. Table 3 provides key information including publication year, dataset size (number of pairs across training, validation, and test splits), domain, language, and public access links. Below, we offer a brief overview of each dataset to contextualize its relevance and usage.
@@ 4.1.1 Conventional Datasets for ATS
LCSTS[103] contains more than 2 million Chinese microblogs from domains including politics, economics, and entertainment, sourced from Sina Weibo’s official accounts such as People’s Daily, the Ministry of National Defense, etc. Summaries are manually annotated, with rules applied to ensure quality (e.g., accounts with over 1 million followers). The dataset is divided into three parts: a large master dataset, a high-quality subset with manual scoring, and a refined test set.
@@
XL-sum[115] is a multilingual abstractive summarization dataset with 1 million article-summary pairs in 44 languages, sourced from the BBC. It focuses on diverse and low-resource languages, offering high-quality, concise summaries. Fine-tuned mT5 models achieve competitive results, making XL-sum the largest dataset for multilingual summarization in both samples and language coverage.
@@
4.1.2 Supplementary and Emerging Datasets
This subsection includes summarization datasets cited fewer than 100 times, encompassing both recently released resources and domain-specific datasets that have not yet achieved widespread adoption. Despite their lower citation counts, these datasets offer valuable diversity in domains, formats, and summarization objectives, making them useful for exploring underrepresented challenges in ATS. Table 4 summarizes their characteristics, including size, domain, and availability.
Table 4: Supplementary datasets with fewer than 100 citations in ATS. The table is formatted in the same manner as the Conventional Dataset Table, exhibiting diversity and serving as a valuable complement to ATS datasets.
@@ 4.1.2 Supplementary and Emerging Datasets
[144] $M$-AVS 2021 2,260 Multimodal English https://github.com/xiyan524/$M$-AVS
@@
As summarized in Tables 3 and 4, current summarization datasets offer several strengths: they include large-scale datasets for training deep neural networks and smaller, refined datasets for evaluation, are predominantly open source and easily accessible, and primarily focus on the news domain. However, there is a gap in high-quality datasets for other domains, limiting research and practical applications in specialized areas such as financial earnings releases [76]. To address this need, we propose methodologies for constructing customized datasets, which are detailed in the following section.
@@ 4.2 Techniques to Build New Datasets
LLM-based Data Generation: LLMs are effective for generating summaries. A straightforward approach is to input original texts along with carefully designed prompts, then fine-tune the model to better grasp the summarization task and produce high-quality outputs (see Section 7). Some studies focus on LLM-based annotation methods. For example, [163] used DialoGPT[164] to transform annotation into tasks like keyword extraction and redundancy detection, improving informativeness and relevance. [165] developed a GPT-3 algorithm for few-shot annotation of medical dialogues, outperforming human-trained models in accuracy and coherence. LLM-based data generation could also be incorporated with human annotators to further improve robustness. For instance, [166] proposed a GPT-3-based scheme combining LLM generation with human evaluation for improved outcomes.
@@ 5.1 Pre-processing methods
Stemming: is the process of converting words with the same root or stem to a basic form by eliminating variable endings like “es” and “ed” [175, 176]. This process was initially employed in open vocabulary text mining, for similar reasons as removing stop words: to reduce computation time and enhance recall in information retrieval [177]. [178] proposed reducing each word to its initial letters, a method referred to as Ultra-stemming, which demonstrated significant improvements.
Sentence Segmentation: splits texts into sentences. The simplest method involves using end markers such as “.”, “?”, or “!”, but it is prone to a lot of interference, including abbreviations like “e.g.” or “i.e.”. To address this issue, simple heuristics and regular expressions are employed [179]. Sentence segmentation is typically necessary for processing long texts [180] or for specific model designs, such as using sentences as nodes in a graph [181].
@@ 5.2 Pre-process Toolkits
The primary tools for English pre-processing with Python are NLTK [^1] https://www.nltk.org/, spaCy [^2] https://spacy.io/ and TextBlob [^3] https://github.com/sloria/TextBlob. NLTK (Natural Language Toolkit) stands out as a leading platform for developing Python programs that handle human language data. It offers a suite of text processing libraries covering classification, tokenization, stemming, tagging, parsing, and semantic reasoning. NLTK also provides wrappers for industrial-strength NLP libraries and maintains an active discussion forum. spaCy, a high-performance NLP library optimized for production use, complements these tools by offering efficient processing pipelines and pre-trained models for English and other languages. Its robust features, such as named entity recognition, dependency parsing, and sentence splitting, make it particularly valuable for tasks like summarization, where structured linguistic analysis is critical for extracting key information. TextBlob, built on NLTK, offers a simplified API for various NLP tasks, including part-of-speech tagging, noun phrase extraction, sentiment analysis, classification, translation, and more. While TextBlob is application-oriented and user-friendly, it sacrifices some flexibility compared to NLTK. For other languages, Jieba [^4] https://github.com/LiveMirror/jieba and HanLP [^5] https://github.com/hankcs/HanLP provide effective solutions for the Chinese language. Jieba is a relatively lightweight tool with its primary function being word segmentation, capable of fulfilling most Chinese word splitting needs. On the other hand, HanLP[184] serves as an NLP toolkit built on PyTorch and TensorFlow, contributing to advancing state-of-the-art deep learning techniques in both academia and industry. It has demonstrated notable results in entity segmentation.
@@
Extractive
summarization
Statistical-based Low resource usage, simple implementation. Prone to irrelevant text and high-scoring unimportant sentences. [185, 186, 187]
@@
Abstractive
summarization
Tree-based Creates organized, concise summaries. May miss contextual relationships in long contexts. [197, 198]
@@
Hybrid
summarization
E2SA Enhances readability, preserves key information High complexity, computation cost. [208, 209, 210]
@@
6 Conventional ATS Methods
As shown in section 2.2, conventional ATS methods are typically classified into extractive, abstractive, and hybrid approaches.
6.1 Extractive Summarization
Extractive summarization selects sentences from the original text through three main steps: calculating sentence importance, sorting sentences based on their importance, and selecting the top-k sentences to form the summary. Sentence importance can be calculated using unsupervised or supervised methods. Extractive ATS based on unsupervised approaches relies on algorithms such as statistical, clustering, and topic-based techniques, offering high performance with low resource requirements, which will be introduced in detail below: